OpenStudy (anonymous):

on a 45 45 90 triangle, what is the value of one leg being y and one leg being blank, hypotenuse 10?

OpenStudy (anonymous):

can you explain how you got that?

jimthompson5910 (jim_thompson5910):

Both legs are equal. If you have a leg of x, then the hypotenuse is \[x\sqrt{2}\] So because the hypotenuse is 10, we know that \[x\sqrt{2}=10\]. Solve for x to get \[x=\frac{10}{\sqrt{2}}=\frac{10\sqrt{2}}{2}=5\sqrt{2}\] So the legs are of length \[5\sqrt{2}\] units